find the horizontal and vertical tangents to

x

=

cos

(

3

t

)

and

y

=

2

sin

(

t

)

?


1
Expert's answer
2021-07-19T10:54:07-0400
"\dfrac{dy}{dx}=\dfrac{\dfrac{dy}{dt}}{\dfrac{dx}{dt}}=-\dfrac{2\cos(t)}{3\sin(3t)}"

Horizontal tangents occur when the derivative equals 0.


"-\dfrac{2\cos(t)}{3\sin(3t)}=0=>t=\dfrac{\pi}{2}+\pi n, n\in \Z"

"x=0, y=\pm2"

Horizontal tangents: "y=-2, y=2."


Vertical tangents occur when the derivative is undefined.


"3\sin(3t)=0=>t=\dfrac{\pi n}{3}, n\in \Z"

"x=\pm1, y=-\dfrac{\pi}{3}, 0, \dfrac{\pi }{3}"

Vertical tangents: "x=-1, x=1."